Please solve this for me by using homogenisation:
2x + 3y – z = 0
x – y – 2z = 0
3x + y + 3z = 0

Grade:11

1 Answers

Vikas TU
7 years ago
Find the determinant first by arranging the termss in a 3x3 matrix form,
Then find Dx by replacing first column of the marix by [0 0 0]T 
similarly find Dy and Dz by replacing the suitable columns respectively.
And put x= Dx/D
y= Dy / D
and z = Dz/D
and D should be non- zero.